# Gregorio Curello Mannheim Ludvig Sinander

### v1

@@ 6 Dominance by objects (rather than mixtures)
For any $\varepsilon\in(0,1/2)$, any non-empty open interval $I\subseteq\mathbf{R}$, and any weakly increasing and non-constant $f:I\to\mathbf{R}$, there exist $u<v$ in $I$ and $\alpha\in[1/2,1/2+\varepsilon]$ such that $\alpha f(v)+(1-\alpha)f(u)>f(\frac{1}{2}u+\frac{1}{2}v)$.
Proof of Lemma˜2.
Suppose toward a contradiction that $f:I\to\mathbf{R}$ is weakly increasing and non-constant and that for some $\varepsilon\in(0,1/2)$, $\alpha f(v)+(1-\alpha)f(u)\leq f\mathopen{}\mathclose{{\left(\frac{1}{2}u+\frac{1}{2}v}}\right)$ for all $u<v$ in $I$ and $\alpha\in[1/2,1/2+\varepsilon]$. Since $f$ is Lebesgue-measurable (being weakly increasing), it follows by Sierpiński’s theorem (section˜2.2) that $f$ is concave. Note that $f$ is differentiable a.e. Choose any $m\in I$ at which $f$ is differentiable, with $f^{\prime}(m)>0$ (which must exist as $f$ is concave, weakly increasing, and non-constant), and observe that
$\lim_{\delta\to 0}\frac{\mathopen{}\mathclose{{\left(1/2+\varepsilon}}\right)f(m+\delta)+\mathopen{}\mathclose{{\left(1/2-\varepsilon}}\right)f(m-\delta)-f(m)}{\delta}=2\varepsilon f^{\prime}(m)>0.$
Hence, we may choose a (small) $\delta>0$ such that
$\mathopen{}\mathclose{{\left(\tfrac{1}{2}+\varepsilon}}\right)f(m+\delta)+\mathopen{}\mathclose{{\left(\tfrac{1}{2}-\varepsilon}}\right)f(m-\delta)>f(m)$
But then $\alpha f(v)+(1-\alpha)f(u)>f\mathopen{}\mathclose{{\left(\frac{1}{2}u+\frac{1}{2}v}}\right)$ holds for $\alpha\coloneqq 1/2+\varepsilon$, $u\coloneqq m-\delta$ and $v\coloneqq m+\delta$, a contradiction. ∎
Consider the following properties, which may or may not be satisfied jointly by a given collection $(I_{s})_{s\in\mathcal{S}}$ of non-empty open intervals of $\mathbf{R}$ and a transformation $\Phi:\mathbf{R}^{\mathcal{S}}\to\mathbf{R}^{\mathcal{S}}$:
  1. (1)
    $\text{NSD}(X,\mathcal{S},U)\subseteq\text{NSD}(X,\mathcal{S},\Phi\circ U)$ for every finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$.
  2. (2)
    $\text{NWD}(X,\mathcal{S},U)\subseteq\text{NWD}(X,\mathcal{S},\Phi\circ U)$ for every finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$.
(The arbitrary intervals $(I_{s})_{s\in\mathcal{S}}$ will not be required in this appendix, where only the special case $I_{s}=\mathbf{R}$ is relevant, but this extra generality will be utilised in the next appendix to prove Proposition˜2.)
Lemma 3.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. For each $s\in\mathcal{S}$, fix a non-empty open interval $I_{s}\subseteq\mathbf{R}$ and a function $\phi_{s}:I_{s}\to\mathbf{R}$, and let $\Phi:\prod_{s\in\mathcal{S}}I_{s}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s}(u_{s}))_{s\in\mathcal{S}}$-situation-wise. Property (1) holds only if either (i) for each $s\in\mathcal{S}$, $\phi_{s}$ is concave and weakly increasing or (ii) for some $s\in\mathcal{S}$, $\phi_{s}$ is constant.
Beweis.
We assume that for all $s\in\mathcal{S}$, $\phi_{s}$ is not constant.
First, suppose that for some $i\in\mathcal{S}$, $\phi_{i}$ is not weakly increasing. Then, there exist $u_{i}>v_{i}$ in $I_{i}$ such that $\phi_{i}(u_{i})<\phi_{i}(v_{i})$. For each $j\in\mathcal{S}\setminus\{i\}$, since $\phi_{j}$ is not constant, there are $u_{j},v_{j}\in I_{j}$ such that $\phi_{j}(u_{j})<\phi_{j}(v_{j})$. This defines a pair of vectors in $\mathbf{R}^{\mathcal{S}}$, $u=(u_{j})_{j\in\mathcal{S}}$ and $v=(v_{j})_{j\in\mathcal{S}}$. Define $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ by $U(x,\cdot)=u$ and $U(y,\cdot)=v$. Note that $U(\mathopen{}\mathclose{{\left\{x,y}}\right\},s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, that $x\in\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},U)$ since $U(x,i)>U(y,i)$, and that $\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},\Phi\circ U)=\mathopen{}\mathclose{{\left\{y}}\right\}$. Hence $\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},U)\nsubseteq\text{NSD}(\mathopen{}\mathclose{{\left\{x,y}}\right\},\mathcal{S},\Phi\circ U)$.
Second, suppose that $\phi_{s}$ is weakly increasing for all $s\in\mathcal{S}$, and that $\phi_{i}$ is not concave for some $i\in\mathcal{S}$. Since $\phi_{i}$ is weakly increasing, it is Lebesgue-measurable, so by Sierpiński’s theorem (section˜2.2), there are $u_{i}<v_{i}$ in $I_{i}$ such that
$\tfrac{1}{2}\phi_{i}(u_{i})+\tfrac{1}{2}\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right).$
Fix an $\varepsilon>0$ small enough that
$\alpha\phi_{i}(u_{i})+(1-\alpha)\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right)$
for all $\alpha\in[1/2,1/2+\varepsilon]$. Fix another $j\in\mathcal{S}\setminus\{i\}$. Since $\phi_{j}$ is non-constant and weakly increasing, by Lemma˜2, there are $u_{j}<v_{j}$ in $I_{j}$ and an $\alpha\in[1/2,1/2+\varepsilon]$ such that
$\alpha\phi_{j}(v_{j})+(1-\alpha)\phi_{j}(u_{j})>\phi_{j}\mathopen{}\mathclose{{\left(\tfrac{1}{2}v_{j}+\tfrac{1}{2}u_{j}}}\right).$
For all $k\in\mathcal{S}\setminus\mathopen{}\mathclose{{\left\{i,j}}\right\}$, there exist $u_{k}<v_{k}$ in $I_{k}$ such that $\phi_{k}(u_{k})<\phi_{k}(v_{k})$ (as each $\phi_{k}$ is non-constant and weakly increasing). So, fix a three-element set $X\coloneqq\{x,y,z\}$, and consider the finite problem $(X,\mathcal{S},U)$ defined by, for all $k\in\mathcal{S}\setminus\mathopen{}\mathclose{{\left\{i,j}}\right\}$,
@@
U(row,column) ijk
Note that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $y\in\text{NSD}(X,\mathcal{S},U)$, and the mixture $p\in\Delta(X)$ given by $p(x)\coloneqq\alpha$, $p(y)\coloneqq 0$ and $p(z)\coloneqq 1-\alpha$ satisfies $\sum_{w\in X}\Phi(U(w,\cdot))p(w)\gg\Phi(U(y,\cdot))$, so that $y\notin\text{NSD}(X,\mathcal{S},\Phi\circ U)$. ∎
Lemma 4.
Assume that $2\leq\mathopen{}\mathclose{{\left\lvert\mathcal{S}}}\right\rvert$$\infty$. For each $s\in\mathcal{S}$, fix a non-empty open interval $I_{s}\subseteq\mathbf{R}$ and a function $\phi_{s}:I_{s}\to\mathbf{R}$, and let $\Phi:\prod_{s\in\mathcal{S}}I_{s}\to\mathbf{R}^{\mathcal{S}}$ be $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise. Property (2) holds only if either (i) for each $s\in\mathcal{S}$, $\phi_{s}$ is concave and strictly increasing or (ii) for every $s\in\mathcal{S}$, $\phi_{s}$ is constant.
Beweis.
We prove the contra-positive. Suppose there is at least one $i^{\star}\in\mathcal{S}$ such that $\phi_{i^{\star}}$ is not constant. We consider three cases:
  1. (1)
    Case 1: for some $i\in\mathcal{S}$, $\phi_{i}$ fails to be weakly increasing.
  2. (2)
    Case 2: case 1 does not hold and, for some $i\in\mathcal{S}$, $\phi_{i}$ fails to be strictly increasing.
  3. (3)
    Case 3: cases 1 and 2 do not hold and, for some $i\in\mathcal{S}$, $\phi_{i}$ fails to be concave.
Case 1: for some $i$, $\phi_{i}$ fails to be weakly increasing. Fix $u<v$ in $I_{i}$ such that $\phi_{i}(u)>\phi_{i}(v)$. Fix a two-element set $X\coloneqq\{x,y\}$, and consider a finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $U(x,i)=u$, $U(y,i)=v$, and $U(x,j)=U(y,j)$ for each $j\in\mathcal{S}\setminus\{i\}$. Note that $\text{NWD}(X,\mathcal{S},U)=\{y\}\nsubseteq\{x\}=\text{NWD}(X,\mathcal{S},\Phi\circ U)$.
Case 2: not case 1, and for some $i$, $\phi_{i}$ fails to be strictly increasing. Recall that there is at least one $i^{\star}\in\mathcal{S}$ whose $\phi_{i^{\star}}$ is not constant.
  • If there exists $k\in\mathcal{S}\setminus\{i\}$ with non-constant $\phi_{k}$, then there exist $u_{i}<v_{i}$ in $I_{i}$ and $u_{k}<v_{k}$ in $I_{k}$ such that $\phi_{i}(u_{i})=\phi_{i}(v_{i})$ and $\phi_{k}(u_{k})<\phi_{k}(v_{k})$.
  • If instead $i$ is the only situation such that $\phi_{i}$ is not constant (i.e., $\phi_{j}$ is constant for every $j\in\mathcal{S}\setminus\{i\}$), then fixing an arbitrary other $j\in\mathcal{S}\setminus\{i\}$, there exist $u_{j}<v_{j}$ in $I_{j}$ and $u_{i}<v_{i}$ in $I_{i}$ such that $\phi_{j}(u_{j})=\phi_{j}(v_{j})$ and $\phi_{i}(u_{i})<\phi_{i}(v_{i})$.
Hence, either way, there exist situations $j\neq k$ in $\mathcal{S}$ and values $u_{j}<v_{j}$ in $I_{j}$ and $u_{k}<v_{k}$ in $I_{k}$ such that $\phi_{j}(u_{j})=\phi_{j}(v_{j})$ and $\phi_{k}(u_{k})<\phi_{k}(v_{k})$. Fix a two-element set $X\coloneqq\{x,y\}$, and consider a finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $U(x,j)=u_{j}$, $U(y,j)=v_{j}$, $U(x,k)=v_{k}$, $U(y,k)=u_{k}$, and $U(x,\ell)=U(y,\ell)$ for each $\ell\in\mathcal{S}\setminus\{j,k\}$. Note that $\text{NWD}(X,\mathcal{S},U)=\{x,y\}\nsubseteq\{x\}=\text{NWD}(X,\mathcal{S},\Phi\circ U)$.
Case 3: not case 1 or case 2, and for some $i$, $\phi_{i}$ fails to be concave. Since $\phi_{i}$ is weakly increasing, it is Lebesgue-measurable, so by Sierpiński’s theorem (section˜2.2), there are $u_{i}<v_{i}$ in $I_{i}$ such that
$\tfrac{1}{2}\phi_{i}(u_{i})+\tfrac{1}{2}\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right).$
Fix an $\varepsilon>0$ small enough that
$\alpha\phi_{i}(u_{i})+(1-\alpha)\phi_{i}(v_{i})>\phi_{i}\mathopen{}\mathclose{{\left(\tfrac{1}{2}u_{i}+\tfrac{1}{2}v_{i}}}\right)$
for all $\alpha\in[1/2,1/2+\varepsilon]$. Fix another situation $j\in\mathcal{S}\setminus\{i\}$. Since $\phi_{j}$ is strictly increasing, by Lemma˜2, there are $u_{j}<v_{j}$ in $I_{j}$ and an $\alpha\in[1/2,1/2+\varepsilon]$ such that
$\alpha\phi_{j}(v_{j})+(1-\alpha)\phi_{j}(u_{j})>\phi_{j}\mathopen{}\mathclose{{\left(\tfrac{1}{2}v_{j}+\tfrac{1}{2}u_{j}}}\right).$
Fix a three-element set $X\coloneqq\{x,y,z\}$, and consider a finite problem $(X,\mathcal{S},U)$ such that $U(X,s)\subseteq I_{s}$ for each $s\in\mathcal{S}$, $U(\cdot,k)$ is constant for all $k\in\mathcal{S}\setminus\{i,j\}$, and

### v2

@@ 6 Dominance by objects (rather than mixtures)
  1. (i)
    If $\Phi$ is not situation-wise, then properties (1) and (2) both fail.
  2. (ii)
    If $\Phi$ is not situation-wise, then properties (3) and (4) both fail.
Proof of Theorem˜1.
Sufficiency follows from Propositions˜1 and 2. For necessity, in light of Propositions˜1 and 2, all that must be shown is that situation-wiseness is necessary for each of properties (1)(4), and this follows from Lemma˜3. ∎
As mentioned in section˜4, Lemma˜3(i) is almost immediate, whereas Lemma˜3(ii) requires a non-trivial proof.
Proof of Lemma˜3(i).
Fix arbitrary $x\neq y$. Suppose that $\Phi$ is not situation-wise; we will show that properties (1) and (2) both fail. Since $\Phi$ is not situation-wise, there exist $s\in\mathcal{S}$ and $u\neq v$ with $u_{s}=v_{s}$ such that $\Phi_{s}(u)<\Phi_{s}(v)$. Then for $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ given by $U(x,\cdot)\coloneqq u$ and $U(y,\cdot)\coloneqq v$, we have $x\in\text{NWD}(\{x,y\},\{s\},U)\subseteq\text{NSD}(\{x,y\},\{s\},U)$ and $x\notin\text{NSD}(\{x,y\},\{s\},\Phi\circ U)\supseteq\text{NWD}(\{x,y\},\{s\},\Phi\circ U)$, so properties (1) and (2) both fail. ∎
Proof of Lemma˜3(ii).
Fix arbitrary $x\neq y\neq z\neq x$. Assume that either property (3) or property (4) holds; we will show that $\Phi$ is situation-wise.
We claim that $\Phi$ is strictly situation-wise monotone in the sense that for any $u,v\in\mathbf{R}^{\mathcal{S}}$ and any situation $s\in\mathcal{S}$, $u_{s}>v_{s}$ implies $\Phi_{s}(u)>\Phi_{s}(v)$. To see why, suppose toward a contradiction that $u_{s}>v_{s}$ and $\Phi_{s}(u)\leq\Phi_{s}(v)$ for some $u,v\in\mathbf{R}^{\mathcal{S}}$ and $s\in\mathcal{S}$. Then letting $U:\{x,y\}\times\mathcal{S}\to\mathbf{R}$ be given by $U(x,\cdot)\coloneqq v$ and $U(y,\cdot)\coloneqq u$, we have $x\notin\text{NSD}(\{x,y\},\{s\},U)\supseteq\text{NWD}(\{x,y\},\{s\},U)$ but $x\in\text{NWD}(\{x,y\},\{s\},\Phi\circ U)\subseteq\text{NSD}(\{x,y\},\{s\},\Phi\circ U)$, so properties (3) and (4) both fail—a contradiction.
For each $s\in\mathcal{S}$, define $\phi_{s}:\mathbf{R}\to\mathbf{R}$ by $\phi_{s}(u_{s})\coloneqq\Phi_{s}(u_{s}\boldsymbol{1}^{s})$ for every $u_{s}\in\mathbf{R}$, where $\boldsymbol{1}^{s}\in\{0,1\}^{\mathcal{S}}$ is given by $\boldsymbol{1}^{s}_{s}\coloneqq 1$ and $\boldsymbol{1}^{s}_{t}\coloneqq 0$ for every $t\in\mathcal{S}\setminus\{s\}$. Strict situation-wise monotonicity implies that for each $s\in\mathcal{S}$, $\phi_{s}$ is strictly increasing and satisfies $\phi_{s}(u_{s}-\varepsilon)<\Phi_{s}(u)<\phi_{s}(u_{s}+\varepsilon)$ for every $u\in\mathbf{R}^{\mathcal{S}}$ and $\varepsilon>0$. Letting $\varepsilon$ vanish yields
$\phi_{s}(u_{s}-)\leq\Phi_{s}(u)\leq\phi_{s}(u_{s}+)\quad\text{for all $u\in\mathbf{R}^{\mathcal{S}}$.}$ ($\ast$)
Hence to show that $\Phi$ is $(\phi_{s})_{s\in\mathcal{S}}$-situation-wise, it suffices to show that $\phi_{s}$ is continuous for each $s\in\mathcal{S}$.
To that end, suppose toward a contradiction that there are $s\in\mathcal{S}$ and $u\in\mathbf{R}$ such that $\phi_{s}(u-)<\phi_{s}(u+)$. It suffices to exhibit $a,b,c\in\mathbf{R}^{\mathcal{S}}$ such that $\text{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})\not\ni c\in\text{NWD}(\{a,b,c\},\mathcal{S},\Phi)$, where $\mathrm{id}(d,\cdot)\coloneqq d$ for each $d\in\{a,b,c\}$, since then properties (3) and (4) both fail—a contradiction.
Fix an arbitrary $t\in\mathcal{S}\setminus\{s\}$. Since $\phi_{s}(u-)<\phi_{s}(u+)$ and since $\phi_{s}$ and $\phi_{t}$ are continuous a.e. (being increasing), we may choose a $v\in(u,\infty)$ such that
$\phi_{s}(u+)>\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}$ ($\dagger$)
and $\phi_{s}$ and $\phi_{t}$ are both continuous at $v$. Since $\phi_{t}$ is continuous at $v$ and strictly increasing, there exists a $w\in(u,v)$ such that
$\phi_{t}(w-)>\frac{\phi_{t}(u+)+\phi_{t}(v)}{2}.$ ($\ddagger$)
Choose $v^{\prime}\in(-\infty,u)$ and $w^{\prime}\in(u,w)$ close enough to $u$ that $\underline{\alpha}\coloneqq(w-v^{\prime})/(v-v^{\prime})$ and $\overline{\alpha}\coloneqq(v-w^{\prime})/(v-v^{\prime})$ satisfy $\underline{\alpha}<\overline{\alpha}$, and such that both $\phi_{s}$ and $\phi_{t}$ are continuous at $v^{\prime}$. Note that $v^{\prime}<u<w^{\prime}<w<v$. Define $a,b,c\in\mathbf{R}^{\mathcal{S}}$ by $a_{s}\coloneqq v$, $b_{s}\coloneqq v^{\prime}$, $c_{s}\coloneqq w^{\prime}$, and $a_{r}\coloneqq v^{\prime}$, $b_{r}\coloneqq v$, $c_{r}\coloneqq w$ for every $r\in\mathcal{S}\setminus\{s\}$. Then, keeping in mind a mixture $p\in\Delta(\{a,b,c\})$ with $\underline{\alpha}<p(b)<\overline{\alpha}$ and p(a)=1p(b)p(a)=1-p(b), we see that $c\notin\text{NSD}(\{a,b,c\},\mathcal{S},\mathrm{id})$.
It remains to show that $c\in\text{NWD}(\{a,b,c\},\mathcal{S},\Phi)$. To this end, note that $b_{s}<c_{s}<a_{s}$ and $a_{t}<c_{t}<b_{t}$, so that $\Phi_{s}(b)<\Phi_{s}(c)<\Phi_{s}(a)$ and $\Phi_{t}(a)<\Phi_{t}(c)<\Phi_{t}(b)$ by strict situation-wise monotonicity. It therefore suffices to show that
@@
For the case $r=s$, we have
$\Phi_{s}(c)\geq\phi_{s}(w^{\prime}-)\geq\phi_{s}(u+)\\ >\frac{\phi_{s}(u-)+\phi_{s}(v)}{2}\geq\frac{\phi_{s}(v^{\prime})+\phi_{s}(v)}{2}=\frac{\Phi_{s}(b)+\Phi_{s}(a)}{2},$